OpenStudy (anonymous):

A small pizza has a diameter of 21 cm. What is the approximate circumference of the pizza? Use pi = 3.14. ( ) cm

OpenStudy (anonymous):

@sissyedgar @ribhu

OpenStudy (sissyedgar):

OMG! i just did this in class earler! i believe its 65.94!

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm d} }\) - diameter\(\large\color{black}{ \displaystyle \\[0.4em] }\) \(\large\color{black}{ \displaystyle {\rm C} }\) - Circumference\(\large\color{black}{ \displaystyle \\[0.4em] }\) There is a formula: \(\large\color{black}{ \displaystyle {\rm C} ={\rm d}\times {\rm \pi} }\) plug in your diameter, \(\large\color{black}{ {\rm d}=21 }\) plug in your approximation for \(\large\color{black}{ {\rm \pi} }\), \(\large\color{black}{ {\rm \pi}=3.14 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm C=~}21~\times~3.14 }\)

OpenStudy (anonymous):

so its 65.94?

OpenStudy (sissyedgar):

like i said!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

65.94

OpenStudy (anonymous):

How I got my answer: Keys: C= Circumference R= Radius D= Diameter 3.14= pi 1) D=21 so I took that and divided it by 2 = 10.5 (our Radius) 2) The formula for the circumference of a circle is 2 x 3.14 x r Plus in the numbers: 2 x 3.14 x 10.5 = 65.94

OpenStudy (anonymous):

*Plug in the numbers*